what is the slope of the line which contains (-3,1 1/2) and passes through the origin?
Answer by padmameesala(42)

the two points on the line are (-3,3/2) and (0,0)
slope = (3/2-0)/(-3-0)
= (3/2)/-3
=-1/2

hope you understand
if any queries youare always welcome